# D2-Mamba: Dual-Scale Fusion and Dual-Path Scanning with SSMs for Shadow Removal

### v2

@@ Introduction
1 - - 37.92 0.980 4.60 39.58 0.981 2.85 34.87 0.955 3.44
2 - - 37.54 0.978 4.72 39.09 0.979 2.94 34.47 0.950 3.53
3 - 38.17 0.981 4.48 39.82 0.982 2.82 35.13 0.956 3.37
4 38.52 0.981 4.37 40.03 0.982 2.76 35.42 0.958 3.30
Table 3: Ablation study of the proposed D2-Mamba on the SRD dataset.
Comparison with State-of-the-Art
We compare the proposed model with popular or state-of-the-art shadow removal methods. Results of all comparison methods are either obtained from their original publications or reproduced using official implementations  [^1] Evaluation metrics for all datasets and experimental results on the ISTD dataset are provided in Appendix..
Quantitative evaluation. Table 1 and Table 2 present the quantitative evaluation results over SRD and ISTD+, respectively. Our method consistently outperforms or matches existing approaches across multiple evaluation metrics, demonstrating its effectiveness and robustness in shadow removal. On the SRD dataset, which contains more complex scenes and textures, diffusion-based methods perform well due to their strong generative capacity for modeling intricate details, significantly outperforming traditional regression-based approaches. However, their long inference time and limited scalability to high-resolution inputs remain key drawbacks. In contrast, our Mamba-based regression model, equipped with accurate cross-scale fusion and similarity-aware grouping, achieves competitive or superior performance with far fewer inference steps, making it both efficient and practical for real-world deployment.
Qualitative evaluation. To further validate visual effectiveness, we provide qualitative comparisons on SRD and ISTD+, which cover diverse scenes with varying shadow shapes, densities, and lighting conditions, providing a comprehensive evaluation of robustness. Figures 3 and  4 clearly show that our method consistently achieves smoother shadow boundaries and more complete removal of residual artifacts compared to other approaches. Some methods produce shadow-recovered regions that appear noticeably brighter or darker than their surroundings(e.g., first and third examples in Figure 3), while others struggle with boundary artifacts (e.g., second and fourth examples). These limitations become more evident in challenging scenarios, such as high-saturation scenes in Figure 4, which often lead to color inconsistencies. In contrast, our approach effectively removes shadow artifacts while restoring natural luminance and color, resulting in improved global consistency and visual realism.
Ablation Study
We conduct ablation studies on SRD to assess the effectiveness of D2-Mamba’s key components. Table 3 shows that each component contributes to performance gains.
@@
Conclusion
@@
Appendix A A. Evaluation Metrics
To enable quantitative comparisons with existing methods, we follow previous studies (Zhu et al. 2022; Fu et al. 2021; Le and Samaras 2021; Wan et al. 2022) and adopt RMSE in the LAB color space to evaluate shadow removal performance. This metric measures the difference between the predicted shadow-free images and the ground truth, where a lower RMSE indicates better restoration quality. In addition, to assess performance in the RGB color space, we also employ PSNR and SSIM as evaluation metrics. All images used for evaluation are uniformly resized to 256×256 to ensure consistency across comparisons.
Appendix B B. Quantitative Results on the ISTD dataset
In this section, we present the detailed quantitative results of shadow removal on the ISTD dataset, comparing our method, D2-Mamba, with several recent state-of-the-art techniques. We compute PSNR, SSIM, and RMSE for the shadow region (S), non-shadow region (NS), and the entire image (ALL).
As shown in Table 4, D2-Mamba demonstrates competitive performance across all these categories. Specifically, in the shadow region (S), our method achieves a PSNR of 39.14, SSIM of 0.992, and RMSE of 5.61, outperforming most state-of-the-art techniques, indicating its ability to effectively remove shadows while preserving image details. In the non-shadow region (NS), D2-Mamba achieves the best results with a PSNR of 34.10, SSIM of 0.981, and a low RMSE of 3.81, highlighting its strength in preserving the quality of the non-shadowed areas. Additionally, in terms of overall image quality (ALL), D2-Mamba maintains a PSNR of 32.18, SSIM of 0.969, and RMSE of 4.06, consistently outperforming many competing methods. These results demonstrate that D2-Mamba offers significant advantages over existing methods, providing excellent shadow removal performance while preserving image quality in both shadow and non-shadow regions. This makes D2-Mamba a strong candidate for real-world shadow removal tasks.
Appendix C C. Qualitative Results on the ISTD dataset
To further validate the effectiveness of the proposed D2-Mamba framework, we present qualitative results on the ISTD dataset in Figure 5. As shown in the visual comparisons, our method produces cleaner shadow-free images with fewer residual artifacts compared to existing approaches. D2-Mamba demonstrates strong performance in preserving texture details and structural consistency across both shadow and non-shadow regions. The removed shadow areas appear visually natural and seamlessly integrated with the surrounding context.
For example, in the first two examples, the first four methods exhibit noticeable artifacts, where the shadow-removed regions fail to blend naturally with the background. In the third image, due to the presence of complex shadow patterns, the compared methods produce results with significant color inconsistencies within the shadowed areas. In the last two images, these methods struggle to restore the true color of the shadow regions, resulting in residual shadows or visible color distortions.
These qualitative results demonstrate that D2-Mamba achieves a more accurate and visually coherent shadow removal, highlighting its strong capacity to model both global semantic structure and local shadow variations.
Appendix D D. Ablation Study on the Path Execution Order in our proposed DPMG module
As shown in Table 5, we investigated the execution order of the two pathways in the DPMG module and reached a key conclusion: the model achieves better performance when the horizontal scanning path precedes the mask-aware adaptive scanning path, compared to the reverse order.
This finding suggests that the horizontal scanning path plays a crucial role as the initial step for global information extraction. A potential explanation is that the horizontal scanning path allows the model to first capture comprehensive contextual information, which provides a solid foundation for the subsequent fine-grained shadow removal. By understanding the overall structure of the image, the model can more effectively distinguish shadowed from non-shadowed regions, leading to more accurate shadow detail recovery during the mask-aware adaptive scanning stage.
In contrast, when the mask-aware adaptive scanning path is executed first, the model may overly focus on local shadow information while missing essential global context. This can result in inconsistencies or distortions when processing complex shadow patterns, ultimately degrading both the precision of shadow removal and structural continuity.
In summary, placing the horizontal scanning path first provides the model with a better global understanding, which in turn allows the subsequent mask-aware adaptive scanning to perform more accurate and structurally coherent shadow removal tasks.
Appendix E E. Extensive Results on the SBU Dataset
Model Param(M) FLOPs(G) Infer(ms) Mem(GB)
ShadowMaskFormer (Li et al. 2025) 2.28 47.79 42 9.29
ShadowDiffusion (Guo et al. 2023b) 55 364.12 70 12.99
HomoFormer (Xiao et al. 2024a) 17.18 71.25 23 7.40
ShadowFormer-L(Guo et al. 2023a) 11.35 129.20 38 6.93
Ours 9.39 40.09 42 11.69
Table 6: Complexity comparison with recent models.
Appendix F F. Complexity Comparison
Table 6 shows the parameter count, computational complexity, inference time, and memory usage. The parameter count (Param(M)) and computational complexity (FLOPs(G)) are calculated using the thop library. Our model (”Ours”) has a relatively low parameter count and good FLOPs, striking a balance between computational complexity and performance, with a short inference time.

### v3

@@ IV-C Real-World Shadow Removal Performance
- - 37.92 0.980 4.60 39.58 0.981 2.85 34.87 0.955 3.44
- - 37.54 0.978 4.72 39.09 0.979 2.94 34.47 0.950 3.53
- 38.17 0.981 4.48 39.82 0.982 2.82 35.13 0.956 3.37
38.52 0.981 4.37 40.03 0.982 2.76 35.42 0.958 3.30
IV-D Complexity Comparison
Table V shows the parameter count, computational complexity, inference time, and memory usage. The parameter count (Param(M)) and computational complexity (FLOPs(G)) are calculated using the thop library. Our model (”Ours”) has a relatively low parameter count and good FLOPs, striking a balance between computational complexity and performance, with a short inference time.
TABLE V: Complexity comparison with recent models.
Model Param(M) FLOPs(G) Infer(ms) Mem(GB)
ShadowMaskFormer [25] 2.28 47.79 42 9.29
ShadowDiffusion [12] 55 364.12 70 12.99
HomoFormer [43] 17.18 71.25 23 7.40
ShadowFormer-L[11] 11.35 129.20 38 6.93
Ours 9.39 40.09 42 11.69
IV-E Ablation Study
We conduct ablation studies on SRD to assess the effectiveness of D2-Mamba’s key components. Table IV shows that each component contributes to performance gains.
@@ IV-E Ablation Study
TABLE VI: Ablation study of the path execution order on the ISTD dataseet.
Method Shadow Region (S) Non-Shadow Region (NS) All Image (ALL)
[…]
@@
Building on the design of DPMG, we further investigate how the execution order of the two pathways affects the overall performance. The results of this ablation study, shown in Table VI, suggest that the model performs better when the horizontal scanning path precedes the mask-aware adaptive scanning path, rather than the reverse order. This finding reveals an important aspect of how these two pathways interact. When the horizontal scanning path is executed first, it helps the model capture comprehensive global information, laying a strong foundation for the more detailed and localized shadow recovery in the subsequent mask-aware adaptive scanning stage. With an understanding of the overall image structure, the model can more effectively distinguish between shadowed and non-shadowed regions, improving the accuracy and precision of shadow removal. This global context facilitates more natural integration of the shadow recovery into the image, preserving the overall visual coherence. On the other hand, if the mask-aware adaptive scanning path is executed first, the model may focus too much on local shadow details while neglecting the broader context. This leads to difficulties in handling complex shadow patterns, resulting in inconsistencies or distortions. The absence of a global structural foundation can cause imprecise shadow removal and disrupt the continuity of the image. Therefore, the ablation results highlight that starting with the horizontal scanning path ensures the model has a better global understanding, which enhances the subsequent mask-aware adaptive scanning and leads to a more accurate and visually coherent shadow removal.
V Conclusion